Q.

A card from a pack of 52 cards is lost. From the remaining 51 cards, n cards are drawn and are found to be spades. If the probability of the lost card to be a spade is 1150, then n is equal to ___________.          [2025]


Ans.

(2)

Let us define the events:

A : Lost card is a spade

B : n cards drawn from 51 cards are spades.

∴  P(A)=1352=14, P(Ac)=1–14=34

   P(B/A)=Cn12Cn51, P(B/Ac)=Cn13Cn51

         P(B)=P(A)P(B/A)+P(Ac)P(B/Ac)

                      =14Cn12Cn51+34Cn13Cn51

∴ P(AB)=P(B/A)P(A)P(B) ⇒ 1150=Cn12Cn51×14P(B)

⇒ 13–n52–n=1150 ⇒ n=2.